OpenStudy (anonymous):

H= 3a + 5b -4c solve for a

OpenStudy (anonymous):

a = (H-5b+4c)/3

OpenStudy (anonymous):

Algebra

OpenStudy (anonymous):

whats the steps in getting that answer? @stoopkid

OpenStudy (anonymous):

1) Subtract all the terms not containing a from both sides: so subtract 5b then subtract -4a (add 4a). 2) You are left with a single term containing a. Divide everything in this term that is not a from both sides, i.e. divide both sides by 3. Let me show you: H = 3a + 5b - 4c 1) The term 5b does not contain a, so subtract it. H - 5b = 3a - 4c 2) The term -4c does not contain a, so subtract it. H - 5b + 4c = 3a 3) Divide everything that is not a in the term 3a. (H - 5b + 4c)/3 = a